### v2

@@ Abstract
Quantum adiabatic optimization has long been expected to outperform classical methods in solving NP-type problems. While this has been proven in certain experiments, its main applications still reside in academic problems where the size of the system to be solved would not represent an obstacle to any modern desktop computer. Here we develop a systematic procedure to find the global optima of the modularity function to discover community structure in complex networks solely relying on pure annealers rather than hybrid solutions. We bypass the one-hot encoding constraints by hierarchically and recursively encoding binary instances of the problem that can be solved without the need to guess the exact penalties for the Lagrange multipliers. We study the variability, and robustness of the annealing process as a function of network size, directness of connections, topology, and the resolution of the communities. We show how our approach produces meaningful and at least equally optimal solutions to state-of-the-art community detection algorithms while maintaining tractable computing times. Lastly, due to its recursive nature, the annealing process returns intermediate subdivisions thus offering interpretable rather than black-box solutions. These dendrograms can be used to unveil normal and pathological hidden hierarchies in brain networks hence opening the door to clinical workflows. Overall, this represents a first step towards an applicable practice-oriented usage of pure quantum annealing potentially bridging two segregated communities in modern science and engineering; that of network science and quantum computing.
@@ Introduction
Real-life systems are composed of multiple individual components whose interactions, albeit generally describable in simple terms, lead to the emergence of non-trivial phenomena. These complex systems can be effectively modeled using complex networks, graphs composed of nodes, and edges or links between them. A large body of work has been addressed to formally characterize the behavior of these networks, focusing on both structural properties and the dynamic processes they support. Among these, unraveling community structure in complex networks has proven to be a crucial problem for the accurate understanding of multiple real-life systems (e.g., neurological disorders). Community structure, as understood by the existence of highly connected groups of nodes, is usually attained by maximizing a modularity function [1] that quantifies the connectivity between intra- and inter-modular nodes in the network [2].
@@
Noteworthy, such function elegantly maps to an infinite-range spin-glass system [3, 4], and is, in fact, a specific instance of a more general class of systems. This evident proximity of the modularity function caused the rapid emergence of methods based on statistical physics [5, 6, 4]. However, the most used solutions are based on mathematical, statistical, and computer science principles [2, 1, 7, 8, 9, 3, 10, 11, 12], where the use of carefully thought heuristics allowed the algorithms to achieve competitive scores and outstanding computing times [13]. Unfortunately, many of these approaches rely, at least partially, on different heuristics which, do not always return network divisions with the highest modularity [14, 15, 16].
Crucially, maximizing the modularity is an NP-hard problem [17], a scenario where quantum annealing (QA) is expected to excel at [18]. QA is grounded on the theoretical result that the final state of the quantum processor will unequivocally be located in the global minima of the energy landscape upon which the annealer is built [19, 20]. This represents a solid argument to deviate from purely algorithmic attempts to find a community structure in complex networks. Unfortunately, one of the main obstacles to QA stems from the fact that qubits are necessarily represented with binary variables of a quadratic cost function. Thus, encoding problems with more than two states (or assignments) requires a specific type of encoding [21, 22]. This, however, demands the addition of constraints assuring the correctness of such an encoding. For example, one-hot encoding is one of the most popular encoding types, but it requires as many constraints as variables so that each element of the processor is associated with one and only one state [23, 24]. The exact penalty for each one of them is added ad-hoc through a set of Lagrange multipliers whose values are unknown and usually need to be tested and fine-tuned for each particular problem. An alternative is to depart from the binary quadratic formulation and use a discrete quadratic model to encode several classes in a $q$-state Potts model type of system [25]. While successful, this approach comes at a huge transparency cost, since the exact calculations are run entirely inside the D-Wave system ecosystem and divided into classical and quantum processes in a manner that is not accessible nor controllable to the user.
@@
In this work, we introduce a novel algorithm that leverages pure quantum annealing (QA) to maximize network modularity and uncover intrinsic community structures. The algorithm employs recursive subdivisions to detect an arbitrary number of communities, eliminating the need for one-hot encoding and its associated constraints. We extend the formalism in [3] to a general case applicable to various types of networks, arbitrary values of the resolution parameter [4], and we show how the function to optimize via QA remains formally independent of all of these. Lastly, we rigorously evaluate the algorithm’s performance across different network topologies and modularity functions, highlighting its strengths compared to current state-of-the-art methods. These results mark a significant step toward the practical application of quantum computing in real-world scenarios, fostering greater integration across scientific disciplines.
Materials and methods
Modularity index and quantum annealing
@@
A graph or network $G=(V,E)$ is a set of vertices or nodes $V$ and edges $E$. Moreover, a community in the graph is a subset of nodes $c\subseteq V$. A community structure is a set of non-overlapping communities $C=\{c_{r}|c_{r}\subseteq V,c_{s}\cap c_{r}=\varnothing,\text{ for }s\neq r\}$. The set of all possible community structures is denoted as $\mathcal{C}$. Importantly, a community is not a subgraph as it doesn’t require the deletion of inter-community edges. Lastly, the quality of a given community structure $C$ can be quantified with the modularity index [7],
$Q=\sum_{c\in C}\left[\frac{L_{c}}{m}-\gamma\left(\frac{d_{c}}{2m}\right)^{2}\right],$ (1)
@@
where $L_{c}$ is the number of links between nodes in the $c$-th community, $d_{c}$ is the sum of the degrees of those nodes, and $m$ is the total number of links (or edges) in the network. Moreover, $\gamma$ is a resolution parameter that effectively accounts for different topological scales [4] at which interactions between nodes happen [26].
@@
A certain community structure is said to be modular if it exhibits a high modularity index when compared to multiple random instances of a given class of random networks (e.g., Erdős-Renyi). Community detection algorithms aim to find a given community structure $C\in\mathcal{C}$ whose associated modularity function $Q$ lies at a global maximum – or its opposite at a minimum,
$\hat{Q}\doteq\max_{\forall C\in\mathcal{C}}\left[Q\right]=-\min_{\forall C\in\mathcal{C}}\left[-Q\right].$ (2)
@@
Even though $\mathcal{C}$ is bounded, the number of possible elements it contains grows faster than exponentially with the size of the network [27]. Furthermore, it has been proven that maximizing modularity is an NP-complete problem [17], thus justifying QA methods [18]. To maximize the modularity index, the first term in Eq. 1 suggests defining clusters containing a large number of edges. On the contrary, the second term aims to minimize the size of such clusters. This is important because any algorithm could be prone to isolate given nodes to artificially boost the modularity index [25].
@@
For a network of two communities, the modularity function in Eq. 1 can be mapped into an Ising system with variables $s_{i}=\{-1,1\}$ indicating that a certain node belongs to either one of the two communities present in the network [3, 4, 24, 25]. For undirected networks,
$Q=\frac{1}{4m}\sum_{ij\in G}B_{ij}s_{i}s_{j}+\frac{(1-\gamma)}{2},$ (3)
@@
are the entries of the modularity matrix, $A_{ij}$ are the entries of the symmetric adjacency matrix, $m=\frac{1}{2}\sum_{ij\in G}A_{ij}$ is the total edge-weight, and $g_{i}=\sum_{j}A_{ij}$ is the node weighted degree. For directed networks [28],
$Q=\frac{1}{2m}\sum_{ij\in G}B_{ij}^{d}s_{i}s_{j}+\frac{(1-\gamma)}{2},$ (5)
@@
Unconstrained hierarchical binary optimization
@@
Let us define a process $\mathcal{P}\doteq(p_{k})_{k=0}^{\infty}$ as a sequence of elements $p_{k}$ such that its series converges to a finite value,
$\hat{Q}^{\mathcal{P}}\doteq\sum_{k=0}^{\infty}p_{k}.$ (7)
@@
Crucially, minimizing $\hat{Q}^{\mathcal{P}}$ need not be straightforward. However, sequentially optimizing each element of the process $p_{k}$ is entirely equivalent,
$\min\left[\hat{Q}^{\mathcal{P}}\right]=\min\left[\sum_{k=0}^{\infty}p_{k}\right]=\sum_{k=0}^{\infty}\min\left[p_{k}\right].$ (8)
@@
Our heuristic is based on the additive nature of $Q$, where each community represents an isolated entity that independently contributes to its value [7]. A natural way to exploit this property is to recursively divide communities into two [3]. Therefore,
$\begin{cases}p_{k}&=-\left[Q(C^{k})-Q(C^{k-1})\right]\\ p_{0}&=-Q(C^{0})\end{cases}$ (9)
where $10\leq|C^{k}|-|C^{k-1}|\leq|C^{k-1}|+$ 1, and $Q(C^{k})$ is the modularity of the network after splitting a given community $c\in C^{k-1}$ into two. It is straightforward to see that $\hat{Q}^{\mathcal{P}}\to Q(C^{\infty})$ with $C^{\infty}$ being a community structure of arbitrary size. Thus, we wish to maximize the modularity by recursively subdividing each community into two.
Formally, there is no guarantee that a given optimized process will univocally yield the partition with the highest modularity index in Eq. 2, $\min\left[\hat{Q}^{\mathcal{P}}\right]\to\hat{Q}$. However, given that the problem itself is NP-complete [17], heuristic processes are needed. Several algorithms have been developed based on the idea that optimizing in smaller steps, rather than all at once, yields competitive results. Among these, the Louvain method is the most popular [10]. In what follows, we derive the Quadratic Unconstrained Binary Optimization (QUBO) cost function for each element of $\mathcal{P}$. We then show how following the approach described here can yield high-quality community structures of arbitrary size without having to use one-hot encoding constraints [23, 24] nor discrete variables [25].
Undirected networks
Following Eq. 1, the modularity of an undirected network of $|C^{k-1}|$ different communities can be decomposed as
$Q(C^{k-1})=\sum_{c\in\tilde{C}^{k-2}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{d_{c}}{2m}\right)^{2}\right]+\frac{L_{k}}{m}-\gamma\left(\frac{d_{k}}{2m}\right)^{2}$
@@
where $\tilde{C}^{k-2}\cup k=C^{k-1}$. Then, the modularity index of the same network after splitting the $k$-th community in two can also be decomposed as
$Q(C^{k})=\sum_{c\in\tilde{C}^{k-2}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{d_{c}}{2m}\right)^{2}\right]+\sum_{r=1}^{2}\left[\frac{L_{k_{r}}}{m}-\gamma\left(\frac{d_{k_{r}}}{2m}\right)^{2}\right]$
where $k_{1}\cup k_{2}=k$. Using Eq. 3 we obtain a compact expression for the elements of $\mathcal{P}$ in terms of the modularity matrix,
$\begin{split}p_{k}&=-\left[Q(C^{k})-Q(C^{k-1})\right]\\ &=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}s_{i}s_{j}-B_{ij}\right)-\frac{1-\gamma}{2}\\ &=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}-\delta_{ij}\sum_{r\in k}B_{ir}\right)s_{i}s_{j}=\frac{-1}{4m}\sum_{ij\in k}B_{ij}^{\text{k}}s_{i}s_{j}\\ &=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j}+\frac{1}{2m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}+\frac{1}{2m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{j}+\frac{1}{4m}\sum_{ij\in k}B_{ij}^{\text{k}}\\ &=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j}.\end{split}$ (10)
We have arbitrarily dropped the constant term $\frac{1-\gamma}{2}$ since it doesn’t play a role in the optimization.
$B_{ij}^{\text{k}}=B_{ij}-\delta_{ij}\sum_{r\in k}B_{ir}\ \ \forall i,j\in k,$ (11)
@@
is the generalized modularity matrix Bk for the $k$-th community [3], we used $s_{i}^{2}=1$, and $\delta_{ij}$ is the Kronecker delta function. In the second row, we have applied a linear mapping from the Ising variables to a set of binary ones $s_{i}=2x_{i}-1$ to obtain the QUBO formulation. The last row follows from some of the properties of the generalized modularity matrix.
@@
Importantly, the first element of the process need not be considered separately. When $k=0$, taking into account that $x_{i}^{2}=x_{i}$, it is easy to show how $p_{k}$ reduces to the known function to split the whole network in two,
$p_{0}=\frac{-1}{m}\left[\sum_{ij\in G}B_{ij}x_{i}x_{j}+(1-\gamma)\sum_{i\in G}g_{i}x_{i}\right]=-Q(C^{0}).$
Even more, it follows that,
$\text{if }x_{i}=x_{j}\ \ \forall i,j\in k\subset V\implies p_{k}=0,$
@@
hence no further increment can be obtained by further subdividing the chosen community. This provides a natural stopping for the hierarchical divisions without the need to predefine a maximum number of communities, either with one-hot encoding dimensions [23, 24] or integer variables [25]. In practical terms, given that the network is symmetric, the summations do need to cover all the $N^{2}$ entries but rather the diagonal and upper diagonal entries (i.e., $\frac{N^{2}}{2}$). This also speeds up the computation parts, where the QUBO problems have significantly shorter symbolic expressions (see Supplementary Material).
Directed networks

### v3

@@ Abstract
Quantum adiabatic optimization has long been expected to outperform classical methods in solving NP-type problems. While this has been proven in certain experiments, its main applications still reside in academic problems where the size of the system to be solved would not represent an obstacle to any modern desktop computer. Here, we develop a systematic procedure to find the global optima of the modularity function to discover community structure in complex networks solely relying on pure annealers rather than hybrid solutions. We bypass the one-hot encoding constraints by hierarchically and recursively encoding binary instances of the problem that can be solved without the need to guess the exact penalties for the Lagrange multipliers. We study the variability and robustness of the annealing process as a function of network size, directness of connections, topology, and the resolution of the communities. We show how our approach produces meaningful and at least equally optimal solutions to state-of-the-art community detection algorithms while maintaining tractable and potentially scalable computing times. Lastly, due to its recursive nature, the annealing process returns intermediate subdivisions, thus offering interpretable rather than black-box solutions. These dendrograms can be used to unveil normal and pathological hidden hierarchies in brain networks, hence opening the door to clinical workflows. Overall, this represents a first step towards an applicable practice-oriented usage of pure quantum annealing, potentially bridging two segregated communities in modern science and engineering: that of network science and quantum computing.
@@ 1 Introduction
Real-life systems are composed of multiple individual components whose interactions, albeit generally describable in simple terms, lead to the emergence of non-trivial phenomena. These complex systems can be effectively modeled using complex networks composed of nodes and edges or links between them. A large body of work has been addressed to formally characterize these networks, focusing on both structural properties and the dynamic processes they support. Unraveling community structure in complex networks has proven to be a crucial problem for the accurate understanding of multiple real-life systems (e.g., neurological disorders). Community structure, as understood by the existence of highly connected groups of nodes, is usually attained by maximizing a modularity function [Newman-Girvan_2004] that quantifies the connectivity between intra- and inter-modular nodes in the network [Girvan2002].
1.1 Related works (I): Algorithms to maximize the modularity
@@
The modularity function elegantly maps to an infinite-range spin-glass system [Newman2006, Reichardt2006]. This caused a rapid emergence of methods based on statistical physics [Massen2005, Reichardt2006], and simulated annealing was used to illustrate the presence of close maxima in the modularity function [Guimera2004]. However, the most common solutions rely on mathematical, statistical, and computer science principles [Girvan2002, Newman-Girvan_2004, Newman2004, Duch2005, Ziv2005, Newman2006, Blondel2008, Traag2019], where the use of carefully thought-out heuristics allowed the algorithms to achieve competitive scores and outstanding computing times [Danon2005]. Yet many of these approaches do not always return community structures with the highest modularity [aref2022, aref2023, Good2010]. In fact, identifying community structure via modularity maximization is not always optimal [Rosvall2008, Peixoto2014, Damle2018, Ghasemian2020], although in certain settings these approaches are formally equivalent to modularity maximization [Bickel2009, Zhao2012, Newman2016]. While we consider these alternative descriptions, we focus on maximizing the modularity.
1.2 Related works (II): Maximizing the modularity with quantum annealing
Crucially, maximizing the modularity is an NP-hard problem [Brandes2008], a scenario where quantum annealing (QA) is expected to excel at [Farhi2001]. QA is grounded in the theoretical principle that, under ideal adiabatic conditions, the quantum processor will converge to the global minimum of the problem’s energy landscape [Johnson2011, Rajak2022]. This represents a solid argument to deviate from purely algorithmic attempts to find a community structure in complex networks.
Unfortunately, one of the main obstacles to QA stems from the fact that qubits are necessarily represented with binary variables of a quadratic cost function. Thus, solving problems with more than two states (or assignments) requires specific encodings and constraints assuring the correctness of the solution [glover_quantum_2022, chancellor_domain_2019]. For example, one-hot encoding, while popular, needs constraints to ensure each variable takes exactly one value from its possible states  [Ushijima-Mwesigwa2017, Negre2020]. Then, the penalty weight is added through Lagrange multipliers, whose values must be empirically tuned for each specific problem  [mayowa_2022].
An alternative is to depart from the binary quadratic formulation and use a discrete quadratic model to encode several classes in a $q$-state Potts system [Wierzbinski2023]. While successful, this approach comes at a huge transparency cost, since the exact calculations are run entirely inside the D-Wave ecosystem and divided into classical and quantum processes in a manner that is not accessible or controllable to the user.
1.3 Summary of contributions
@@
In this work, we introduce a novel algorithm that leverages pure QA to maximize network modularity and uncover intrinsic community structures. The algorithm employs recursive subdivisions to detect an arbitrary number of communities, eliminating the need for one-hot encoding and its associated constraints. We extend the formalism in [Newman2006] to a general case applicable to various types of networks, arbitrary values of the resolution parameter [Reichardt2006], and we show how the function to optimize via QA remains formally independent of all of these. Lastly, we rigorously evaluate the algorithm’s performance and scalability across different network topologies and modularity functions, highlighting its strengths compared to current state-of-the-art methods. These results mark a significant step toward the practical application of quantum computing in real-world scenarios, fostering greater integration across scientific disciplines.
2 Materials and methods
2.1 Community structure and modularity index
@@
A graph or network $G=(V,E)$ is a set of vertices or nodes $V$ and edges $E$. Moreover, a community in the graph is a subset of nodes $c\subseteq V$. A community structure is a set of non-overlapping communities $C=\{c_{r}|c_{r}\subseteq V,c_{s}\cap c_{r}=\varnothing,\text{ for }s\neq r\}$. The set of all possible community structures is denoted as $\mathcal{C}$. Importantly, a community is not a subgraph as it doesn’t require the deletion of inter-community edges. Lastly, the quality of a given community structure $C$ can be quantified with the modularity index [Newman2004],
$Q(C)=\sum_{c\in C}\left[\frac{L_{c}}{m}-\gamma\left(\frac{d_{c}}{2m}\right)^{2}\right],$ (1)
@@
where $L_{c}$ is the number of links between nodes in the $c$-th community, $d_{c}$ is the sum of the degrees of those nodes, and $m$ is the total number of links (or edges) in the network. Moreover, $\gamma$ is a resolution parameter that effectively accounts for different topological scales [Reichardt2006] at which interactions between nodes happen [Delvenne2010]. It also represents a resolution limit, where communities smaller than a given size cannot be detected by maximizing the modularity  [Fortunato2007, Kumpula2007]. A certain community structure is said to be modular if it exhibits a high modularity index when compared to multiple random instances of a given class of random networks (e.g., Erdős-Renyi).
@@
Community detection algorithms aim to find a given community structure $C\in\mathcal{C}$ whose associated modularity function $Q$ lies at a global maximum – or its opposite at a minimum,
$\hat{Q}\doteq\max_{\forall C\in\mathcal{C}}\left[Q(C)\right]=-\min_{\forall C\in\mathcal{C}}\left[-Q(C)\right].$ (2)
@@
More precisely, $\hat{Q}$ is the modularity of the network, since it is obtained from the community structure that maximizes $Q$ in Eq. 1. To maximize the modularity index, the first term in Eq. 1 suggests defining clusters containing a large number of edges. On the contrary, the second term aims to minimize the size of such clusters. This is important because any algorithm could be prone to isolating given nodes to boost the modularity index [Wierzbinski2023]. Even though $\mathcal{C}$ is bounded, the number of possible elements it contains grows faster than exponentially with the size of the network [Fortunato2007].
@@
Hereon, we denote a community structure of size $|C|=k$ as $C^{k}$. For a network of two communities (i.e., $k=2$), the modularity function in Eq. 1 can be mapped into an Ising system with variables $s_{i}=\{-1,1\}$ indicating that a certain node belongs to either one of the two communities present in the network [Newman2006, Reichardt2006]. For undirected networks, applying the transformation to Eq. 1 yields
$Q(C^{2})=\frac{1}{4m}\sum_{ij\in G}B_{ij}s_{i}s_{j}+\frac{(1-\gamma)}{2},$ (3)
@@
are the entries of the modularity matrix, $A_{ij}$ are the entries of the symmetric adjacency matrix, $m=\frac{1}{2}\sum_{ij\in G}A_{ij}$ is the total edge-weight, and $g_{i}=\sum_{j}A_{ij}$ is the node weighted degree. For directed networks [Leicht2008], applying the transformation to Eq. 1 yields
$Q(C^{2})=\frac{1}{2m}\sum_{ij\in G}B_{ij}^{d}s_{i}s_{j}+\frac{(1-\gamma)}{2},$ (5)
@@
2.2 Unconstrained hierarchical binary optimization
@@
To exploit QA without considering discrete variables and encodings, we wish to maximize the modularity by recursively subdividing each community into two. For that, let us define a process $\mathcal{P}\doteq(p_{k})_{k=0}^{K}$ as a sequence of elements $p_{k}$ such that its series converges to a finite value,
$\hat{Q}^{\mathcal{P}}\doteq\sum_{k=1}^{K}p_{k},$ (7)
@@
where $K$ is any positive integer indicating the number of detected communities. Minimizing $\hat{Q}^{\mathcal{P}}$ need not be straightforward, but optimizing each element is equivalent,
$\min\left[\hat{Q}^{\mathcal{P}}\right]=\min\left[\sum_{k=1}^{K}p_{k}\right]=\sum_{k=1}^{K}\min\left[p_{k}\right].$ (8)
@@
Our heuristic is based on the additive nature of $Q$, where each community represents an isolated entity that independently contributes to its value [Newman2004]. A natural way to exploit this property is to consider the increments in the modularity after recursive divisions [Newman2006]. Therefore,
$p_{k}=-\left[Q(C^{k+1})-Q(C^{k})\right]$ (9)
[…]
Formally, there is no guarantee that a given optimized process will univocally yield the partition with the highest modularity index in Eq. 2, i.e., $\min\left[\hat{Q}^{\mathcal{P}}\right]\to\hat{Q}$. In what follows, we derive the Quadratic Unconstrained Binary Optimization (QUBO) cost function for each element of $\mathcal{P}$.
2.3 Undirected networks
From Eq. 1, the modularity of an undirected network with $k$ communities can be decomposed as
$Q(C^{k})=\sum_{c\in C^{k-1}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{d_{c}}{2m}\right)^{2}\right]+\frac{L_{k}}{m}-\gamma\left(\frac{d_{k}}{2m}\right)^{2}.$
@@ 2.3 Undirected networks
Then, the modularity index of the same network after splitting the $k$-th community in two can also be decomposed as
$Q(C^{k+1})$ $=\sum_{c\in C^{k-1}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{d_{c}}{2m}\right)^{2}\right]$
$\quad\quad\quad\quad+\sum_{r=1}^{2}\left[\frac{L_{k_{r}}}{m}-\gamma\left(\frac{d_{k_{r}}}{2m}\right)^{2}\right],$
where $k_{1}\cup k_{2}=k$. Using Eq. 3 we obtain a compact expression for the elements of $\mathcal{P}$,
$\begin{split}p_{k}&=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}s_{i}s_{j}-B_{ij}\right)\\ &=\frac{-1}{4m}\sum_{ij\in k}\left(B_{ij}-\delta_{ij}\sum_{r\in k}B_{ir}\right)s_{i}s_{j}\\ &=\frac{-1}{4m}\sum_{ij\in k}B_{ij}^{\text{k}}s_{i}s_{j}\\ &=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j},\end{split}$ (10)
where we used $s_{i}^{2}=1$, the Kronecker delta function $\delta_{ij}$, and Bkk is the generalized modularity matrix for the $k$-th community [Newman2006],
$B_{ij}^{\text{k}}=B_{ij}-\delta_{ij}\sum_{r\in k}B_{ir}\ \ \forall i,j\in k.$ (11)
@@
In the fourth row, we have applied a linear mapping from the Ising variables to a set of binary ones $s_{i}=2x_{i}-1$ to obtain the QUBO formulation. The last row follows from properties of the generalized modularity matrix(Supplementary Material S1).
@@
Importantly, the first element of the process need not be considered separately. When $k=1$, taking into account that $x_{i}^{2}=x_{i}$, $p_{k}$ reduces to the function to split the whole network in two,
$\frac{-1}{m}\left[\sum_{ij\in G}B_{ij}x_{i}x_{j}+(1-\gamma)\sum_{i\in G}g_{i}x_{i}\right]=-Q(C^{2})+1-\gamma.$
The minimum of $p_{1}$ is independent of the terms $1-\gamma$ and can be arbitrarily dropped; we only wrote them for completeness. Even more,
$\text{if }x_{i}=x_{j}\ \ \forall i,j\in k\implies p_{k}=0,$
@@
which provides a natural stopping criterion for the hierarchical divisions without the need to predefine a maximum number of communities [Negre2020, Wierzbinski2023]. In practical terms, given that the network is symmetric, the summations don’t need to cover all the $N^{2}$ entries but rather the diagonal and upper diagonal entries.
2.4 Directed networks